OpenStudy (anonymous):

solve for h: h+3/8 = -1/4. will give metal to best answer and become a fan.

jigglypuff314 (jigglypuff314):

subtract 3/8 from both sides so h = -2/8 - 3/8 =

OpenStudy (anonymous):

5/8th?

jigglypuff314 (jigglypuff314):

negative but yeah :)

OpenStudy (anonymous):

ok, thanks!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

can you help me with a couple more? @jigglypuff314

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

find the quotient. simplify if possible: 2 2/3 / 1/4

OpenStudy (anonymous):

\[2\frac{ 2 }{ 3 } \times \frac{ 1 }{ 4 }\]

jigglypuff314 (jigglypuff314):

did you mean \[2\frac{ 2 }{ 3 } \div \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

oh, right. lol

jigglypuff314 (jigglypuff314):

which equals \[2\frac{ 2 }{ 3 } \times 4\]

OpenStudy (anonymous):

ok thanks! i have 2 more if that is ok?

OpenStudy (anonymous):

Go Jiggly, ur a good answerer

OpenStudy (anonymous):

wooop wooop

jigglypuff314 (jigglypuff314):

yeah, sure harly :)

jigglypuff314 (jigglypuff314):

and thanks helpmehplz ;)

OpenStudy (anonymous):

solve for t: 15 3/4 = t+4 5/8

jigglypuff314 (jigglypuff314):

subtract 4 5/8 from both sides so \[t = 15\frac{ 6 }{ 8 } - 4\frac{ 5 }{ 8 }\]

OpenStudy (anonymous):

ok, one sec, let me figure this out. lol

OpenStudy (anonymous):

11 2/8th?

jigglypuff314 (jigglypuff314):

close ;) 6/8 - 5/8 = 1/8 so you get 11 1/8

OpenStudy (anonymous):

ok, thanks. last one. you are so awesome! solve for h: 3/7 h = 9

jigglypuff314 (jigglypuff314):

multiply both sides by 7/3 so h = 9 * 7/3 =

OpenStudy (anonymous):

6.6?

jigglypuff314 (jigglypuff314):

ummm... 9 * 7 = 63 then 63 / 3 =

OpenStudy (anonymous):

lol, i am not all to sure what to do

jigglypuff314 (jigglypuff314):

h = 9 * 7/3 = 9 times 7 is 63 right? then what is 63 divided by 3 =?

OpenStudy (anonymous):

21

jigglypuff314 (jigglypuff314):

yep! :)

OpenStudy (anonymous):

great! thanks so much!

jigglypuff314 (jigglypuff314):

you're welcome ;)